# Censored extreme value estimation

### v1

@@ 3 Extreme Kaplan–Meier Integrals
Weak consistency of extreme Kaplan–Meier integrals under weaker assumptions than Condition 1, may be possible with an approach as in Theorem 1.1 in [39], where the existence of the first moment is assumed for consistency (to the possibly modified target cdf) of the standard Kaplan–Meier integral. However, the proof of the latter result uses some specialized techniques which fall far from our setup, in the like of reverse-time supermartingales. In order not to further overload the notation and proofs, and to keep the narrative coherent, we simply continue to use the natural Condition 1 in Theorem 3.4.
The next theorem states the asymptotic normality of EKM integrals under the weak assumptions of Theorem 3.3. However, note that the centering sequence is random.
@@
The variance of the random variable $W^{\circ}(\varphi)$ turns out to be the correct asymptotic variance in all of our weak convergence results.
Example 1.
Consider $\varphi(x)=\log(x)$. Then
$\gamma_{0}^{\circ}(V^{\circ})$ $=(V^{\circ})^{1/\gamma_{G}},\quad\gamma_{1}^{\circ}(V^{\circ})=(V^{\circ})^{1/\gamma_{G}}(\log(V^{\circ})+\gamma_{F}),$
$\gamma_{2}^{\circ}(V^{\circ})$ $=(V^{\circ})^{1/\gamma_{G}}(\log(V^{\circ})+\gamma_{F}-\gamma_{G})+\gamma_{G}-\gamma_{F},$
and so
$W^{\circ}(\varphi)=(V^{\circ})^{1/\gamma_{G}}(\gamma_{G}-\delta^{\circ}\gamma_{F})+\gamma_{F}-\gamma_{G}.$
Similarly, for $\varphi(x)=x$ we get
$W^{\circ}(\varphi)=(V^{\circ})^{1+1/\gamma_{G}}\Big{(}\delta^{\circ}+\frac{1-\delta^{\circ}}{1-\gamma_{F}}-\frac{1}{(1-\gamma_{F})(1+\gamma_{G})}\Big{)},$
and for $\varphi(x)=1\{x\leq x_{0}\}$ we obtain
$W^{\circ}(\varphi)$ $=1\{V^{\circ}\leq x_{0}\}\Big{(}1+\frac{\gamma_{H}}{\gamma_{G}}x_{0}^{-1/\gamma_{F}}((V^{\circ})^{1/\gamma_{H}}-1)-(V^{\circ})^{1/\gamma_{H}}x_{0}^{-1/\gamma_{F}}(1-\delta^{\circ})\Big{)}$
$\>\>+1\{V^{\circ}>x_{0}\}\Big{(}x_{0}^{1/\gamma_{G}}\frac{\gamma_{H}}{\gamma_{F}}+x^{-1/\gamma_{F}}\frac{\gamma_{H}}{\gamma_{G}}-1\Big{)}.$
@@
In the next result, we replace in (3.9) the random centering sequence $\int\varphi\mathop{}\!\mathrm{d}F^{Z_{n-k,n}}$ with its nonrandom counterpart $\int\varphi\mathop{}\!\mathrm{d}F^{U_{H}(n/k)}$, where $U_{H}$ is the tail quantile function of $H,$ i.e.
$U_{H}(x)=\inf\{y:H(y)\geq 1-1/x\},\quad x>1.$
For this purpose, we need to impose the following condition on the tail quantile functions $U_{F}$ and $U_{H}:$ let a tail quantile function $U$ possess a derivative $U^{\prime}$ and let the following von Mises’ convergence hold
$\lim_{t\to\infty}\frac{tU^{\prime}(t)}{U(t)}=\gamma,$ (3.10)
which in itself is sufficient for $U$ to be regularly varying with tail index $\gamma$, and is standard in extreme value theory (see e.g. Theorem 1.1.11 and Corollary 1.1.12 in [15]). Notice that the condition (3.10) is equivalent to the following
$\lim_{t\to\infty}\frac{tK^{\prime}(t)}{1-K(t)}=\frac{1}{\gamma},$ (3.11)
where $U$ is the tail quantile function of cdf $K.$ Also, we need the following technical condition.
@@
Let $\varphi$ have a derivative $\varphi^{\prime}$ such that for some $\varepsilon^{\prime}>0$
$\int_{1}^{\infty}\bar{\varphi^{\prime}}(x)\,x^{-1/\gamma_{F}+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}x<\infty.$ (3.12)
Hereinafter $\bar{\varphi^{\prime}}$ denotes an envelope of $\varphi^{\prime}.$ If the behaviour of $\varphi^{\prime}$ is regular enough, in particular, if $\varphi^{\prime}$ is continuous, then Condition 2 does not pose any additional complications, since it then follows from Condition 1.
Theorem 3.7.
Assume the conditions of Theorem 3.3. Furthermore, assume $U_{F}$ and $U_{H}$ satisfy (3.10), and $\varphi$ satisfies Condition 2. Then
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(F^{Z_{n-k,n}}-F^{U_{H}(n/k)})\stackrel{{\scriptstyle P}}{{\to}}0.$ (3.13)
As an immediate consequence of Theorems 3.6 and 3.7 we get the following result.
Corollary 3.8.
Under the assumptions of Theorem 3.7, we have
$\sqrt{k}\int\varphi\,\mathop{}\!\mathrm{d}(\amsmathbb{F}_{k,n}-F^{U_{H}(n/k)})\stackrel{{\scriptstyle d}}{{\to}}N(0,\sigma^{2}_{\varphi}).$ (3.14)
The convergence result of Corollary 3.8 is of independent interest, since it is not always possible to replace a nonrandom centering sequence with a constant in limit relations appearing in extreme value theory. For instance, it is not possible in Smirnov’s lemma and its corollary regarding limit behavior of intermediate order statistics, see e.g. Theorems 2.2.1 and 2.4.1 in [15].
Nevertheless, such a non-random replacement is possible in our case if the intermediate sequence $k=k(n)$ does not grow too fast with respect to $n.$ The precise speed is determined through the quality of the regular variation property of $U,$ which in turn is governed by a second order parameter denoted by $\rho,$ see [33]. Thus, to control this speed it is necessary to impose the following standard second order condition.
Condition 3.
@@
Assume the conditions of Theorem 3.3. Furthermore, assume $\varphi$ satisfies Condition 2 and $U_{F}$ satisfies the second order Condition 3. Let $\sqrt{k}\,a(c_{n})\to\lambda$ with finite $\lambda$ and $c_{n}=(1-F(U_{H}(n/k)))^{-1},$ and if $\lambda\neq 0$ assume additionally that $\varphi^{\prime}$ is a.e. continuous. Then
@@
As an immediate consequence of Theorems 3.6 and 3.9, we get the following result.
@@
Notice that for Theorem 3.9 to hold, we have required Condition 1. In particular, if $\varphi$ grows slower than a power function, the condition $\gamma_{F}<\gamma_{G}$ implies the required integrability. This non-heavy censoring regime for asymptotic normality of censored extremes has been observed before for related estimators (related to $\varphi=\log$), see for instance Remark 1 in [6].
@@
Analysing the joint behaviour of empirical versions of $H^{0,t}$ and $H^{1,t}$ is a somewhat direct route to prove asymptotic normality of $S_{k,n}(\varphi)$ using the functional delta method. However, there is an important caveat in that such assertion can only be proved for differentiable $\varphi$ and integrals up to an endpoint $T$ with $F(T)<1$, which is equivalent to $\varphi$ satisfying Condition 1. The above approach allows us to establish asymptotic results on the maximal interval, which in turn allows integration of infinite-support $\varphi$ functions.
4 Residual estimators revisited
4.1 Generalized residual estimators
As before, let $X_{1},\dots,X_{n}$ denote an i.i.d. sequence of random variables with cdf $F,$ and $X_{1,n}\leq\ldots\leq X_{n,n}$ denote the order statistics of $\{X_{i}\}_{i=1}^{n}.$ Denote by $U_{F}(t),$ the tail quantile function of $F,$ i.e. $U_{F}(t)=\inf\{x:F(x)\leq 1-1/t\}.$ For when $F$ belongs to the Fréchet max-domain of attraction (write $F\in\mathcal{D}(G_{\gamma})$ with $\gamma>0$), [34] studied the limit behavior of statistics of the form
$R_{k,n}=\frac{1}{k}\sum_{i=1}^{k}\vartheta(X_{n-i+1,n}/X_{n-k,n}),$
where $\vartheta:[1,\infty)\to\amsmathbb{R}$ is some suitable function, which is here assumed to be almost everywhere continuous.
Theorem 4.1 (Theorem 4.1, [34]).
@@ 4.1 Generalized residual estimators
In further theorems 4.2 and 4.5 ibid., it was proven that $\mu(\xi_{n-k,n})$ in (4.1) can be replaced with the nonrandom limit $\amsmathbb{E}[\vartheta(\xi^{\gamma})]$ under some additional conditions imposed on $F,$ $\vartheta$ and $k.$
It turns out that the direct extension of Theorem 4.1 to the Gumbel and Weibull max-domains of attraction leads to a degenerate limit in (4.1) in almost all cases. Recall the notation $\tau_{F}=\inf\{x:F(x)=1\},$ the right endpoint of $F.$ Specifically, we have the following result, whose proof is analogous to the proof of Theorem 4.1, and thus omitted.
Theorem 4.2.
@@
A reason of a degenerate limit in (4.2) is that for $\gamma\leq 0$ and $\tau_{F}>0$ the ratio $U_{F}(xt)/U_{F}(t)$ tends to $1$ as $t\to\infty$ (which follows from techniques in [15]), see for details Theorem 4.4 and Corollary 4.6.
It follows that to derive a non-degenerate limit, a faster than $\sqrt{k}$ normalizing sequence is required in (4.2). Consequently, a downside of using a residual estimators technique to develop extreme value estimators (in presence of random censorship) is that we may be faced with situations where positive and negative values of the extreme value index are allowed, but the rate of convergence of the estimator inconveniently varies. Another disadvantage is that $\vartheta(1)$ does not depend on $\gamma<0.$ This makes it impossible to construct extreme value index estimators based solely on $R_{k,n}.$
For this reason, let us consider an alternative and general approach. As before, let $\xi,\xi_{1},\xi_{2},\ldots$ be independent Pareto$(1)$ random variables with $n$-th order statistics denoted by $\xi_{1,n}\leq\ldots\leq\xi_{n,n}.$
@@
and for all $x$ there exists $T>0$, such that for $t\geq T$
@@
Let $\vartheta:[0,\infty)\to\amsmathbb{R}$. We are interested in the limit behavior of the following statistic
$R^{f}_{k,n}=\frac{1}{k}\sum_{j=1}^{k}\vartheta\left(f\Big{(}\frac{\xi_{n-j+1,n}}{\xi_{n-k,n}},\xi_{n-k,n}\Big{)}\right),$
where $k=k_{n}$ is an intermediate sequence.
Remark 4.3.
For technical reasons we generally assume that $g_{1},\,g_{2}$ can be modified to tend to $\infty$ together with $f(x,t)$. However, if $\bar{\vartheta}$ and $\bar{\vartheta^{\prime}}$ are eventually non-decreasing (resp. non-increasing), then it suffices to require $g_{2}$ (resp. $g_{1}$) to tend to $\infty$ together with $f(x,t).$
@@
Let $k=k_{n}$ be an intermediate sequence, and assume for some $\delta>0$
$\amsmathbb{E}\Big{[}\max\big{\{}\bar{\vartheta}(g_{1}(\xi)),\bar{\vartheta}(g_{2}(\xi))\big{\}}\Big{]}^{1+\delta}<\infty.$ (4.6)
Then $R^{f}_{k,n}$ is a consistent estimator for $\amsmathbb{E}[\vartheta(g(\xi))].$ If, moreover, (4.6) is satisfied for some $\delta>1,$ then
@@
To replace $\mu_{f}(\xi_{n-k,n})$ with $\mu_{f}(\infty):=\amsmathbb{E}[\vartheta(g(\xi))]$ in Theorem 4.4, we need to impose another condition (compare with Condition 3).
Condition 4.
@@
Let the conditions of Theorem 4.4 and Condition 4 hold. Assume $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\lambda$ with finite $\lambda,$ $\vartheta$ has a derivative ϑ,superscriptitalic-$\vartheta^{\prime},$ and ϑsuperscriptitalic-$\vartheta^{\prime}$ is almost everywhere continuous if $\lambda\neq 0.$ Assume also
$\amsmathbb{E}\Big{[}\hat{g}(\xi)\max\big{\{}\bar{\vartheta^{\prime}}(g_{1}(\xi)),\bar{\vartheta^{\prime}}(g_{2}(\xi))\big{\}}\Big{]}<\infty.$ (4.10)

### v5

@@ 3 Extreme Kaplan–Meier Integrals
Weak consistency of extreme Kaplan–Meier integrals under laxer assumptions than Condition 1 may be possible to establish with an approach as in Theorem 1.1 in [33], where the existence of the first moment of the (possibly modified) target cdf is assumed for consistency of the standard Kaplan–Meier integral. The proof of the latter result uses some specialized techniques which fall far from our setup, for instance by using reverse-time supermartingales.
The next theorem states the asymptotic normality of EKM integrals under the assumptions of Theorem 3.1. However, note that the centering sequence is random.
@@
The variance of $W^{\circ}(\varphi)$ turns out to be the correct asymptotic variance in all of our weak convergence results.
@@
Remark 3.5.
It is possible to replace the random centering sequence $\int\varphi\mathop{}\!\mathrm{d}F^{Z_{n-k,n}}$ in (3.9) with its non-random counterpart $\int\varphi\mathop{}\!\mathrm{d}F^{U_{H}(n/k)}$, where $U_{H}$ is the tail quantile function of $H,$ i.e.
$U_{H}(x)=\inf\{y:H(y)\geq 1-1/x\},\quad x>1,$
under a mild condition imposed on $U_{H}$ and $U_{F}.$ We provide this result in Section B of the Supplementary Material. This is of independent interest, since it is not always possible to replace a non-random centering sequence with a constant in limit relations appearing in extreme value theory.
A non-random replacement of the centering sequence is possible in our case if the intermediate sequence $k=k(n)$ does not grow too fast with respect to $n,$ as is precised in Theorem 3.6 below. The exact speed is determined through the quality of the regular variation property of $U_{F},$ which in turn is governed by a second-order parameter denoted by $\rho,$ see [27]. Thus, to control this speed it is necessary to impose the following standard second-order condition.
@@
Assume the conditions of Theorem 3.1. Furthermore, assume $\varphi$ has a continuous and eventually monotone derivative $\varphi^{\prime},$ and $U_{F}$ satisfies the second-order Condition 2. Let $\sqrt{k}\,a(c_{n})\to\lambda$ with finite $\lambda$ and $c_{n}=(1-F(U_{H}(n/k)))^{-1}.$ Then
@@
As an immediate consequence of Theorems 3.4 and 3.6, we get the following result.
@@
Notice that for Theorem 3.6 to hold, we have required Condition 1. In particular, if $\varphi$ grows slower than a power function, the condition $\gamma_{F}<\gamma_{G}$ implies the required integrability. This non-heavy censoring regime for asymptotic normality of censored extremes has been observed before for related estimators; see for instance Remark 1 in [6].
As an illustration of Theorem 3.6 and Corollary 3.7, let us consider asymptotic behavior of the censored analog of the celebrated Hill estimator [22] of a positive extreme value index. Setting $\varphi=\log$ in (3.3), we get
$\widehat{\gamma}^{c,Hill}_{k}=S_{k,n}(\log)=\int_{1}^{\infty}\log(x)\amsmathbb{F}_{k,n}({\rm d}x).$ (3.14)
This estimator was introduced in equation (7) of [35], under a very different but equivalent expression, and a consistency result was provided. To the best of our knowledge, its asymptotic normality has not been established yet. The following result fills this gap in the literature.
Corollary 3.9.
Let $F$ and $G$ be regularly varying and eventually continuous cdfs with $\gamma_{G}>\gamma_{F},$ Condition 2 hold for $U_{F}$ and $k=k_{n}$ be an intermediate sequence. Let $\sqrt{k}\,a(c_{n})\to\lambda$ with finite $\lambda$ and $c_{n}=(1-F(U_{H}(n/k)))^{-1}.$ Then
[…]
@@
Notice that $S_{k,n}(1)=\int 1\mathop{}\!\mathrm{d}\amsmathbb{F}_{k,n}=\amsmathbb{F}_{k,n}(\infty)\leq 1$, and then $\amsmathbb{F}_{k,n}/S_{k,n}(1)$ is a true cdf. One obtains from Corollary 3.7 that $\sqrt{k}(S_{k,n}(1)-1)\stackrel{{\scriptstyle P}}{{\to}}0$. In particular, this means that estimators normalized by $\amsmathbb{F}_{k,n}(\infty)$ have the same distributional limits as their non-normalized counterparts. Consequently,
$\widehat{\gamma}^{c,Hill}_{k}/\amsmathbb{F}_{k,n}(\infty)-\widehat{\gamma}^{c,Hill}_{k}=o_{\amsmathbb{P}}(k^{-1/2}).$
The above relation is practically important, since normalized estimators can have slightly better finite-sample performance. We adopt normalization in the simulation studies of Section 7.
It is also of independent interest to investigate the asymptotic properties of the extreme Kaplan–Meier estimator $\amsmathbb{F}_{k,n}$ itself. Consider a function $\varphi_{0}(x)=I(x\leq x_{0})$ for some $x_{0}>1,$ then $S_{k,n}(\varphi_{0})=\amsmathbb{F}_{k,n}(x_{0}).$ Since $\varphi_{0}$ is not continuous, one can consider a EKM integral of a continuously differentiable envelope function $\varphi_{\epsilon},$ and then taking $\epsilon\to 0$ for $S_{k,n}(\varphi_{\epsilon})$.
Corollary 3.11.
Let $F$ and $G$ be regularly varying and eventually continuous cdfs with $\gamma_{G}>\gamma_{F}.$ Then $\amsmathbb{F}_{k,n}(x_{0})$ is consistent for $F^{\circ}(x_{0}).$ Moreover, under the assumptions of Corollary 3.9 we have
[…]
In particular, the above result can be used to prove consistency of censored analogues of probability weighted moments and maximum likelihood methods, which are well accepted in extreme value theory, see, e.g., Chapters 3.4 and 3.6.1 in [11], and often applied to extreme quantile and tail estimation.
4 Generalized residual estimators for non-censored data
Thus far we have studied EKM integrals when distributions of both the event and the censoring variables are heavy-tailed. However, the extension of this methodology to all max-domains of attraction is not straightforward, as we precise below. In particular, the limits in probability of EKM integrals may not depend on $\gamma_{F}$ if the event variable distribution is not heavy-tailed, making them unsuitable for estimation.
To overcome this difficulty and find an appropriate normalization of EKM integrals leading to its non-degenerate limit behavior, we turn in this section to the theory of residual estimators developed in [28] for heavy-tailed distributions and extend it to all max-domains of attraction. This route does not suffer normalization drawbacks. A key role in this extension is played by the so-called candidate functions having certain regular limit behavior; see Chapter 1 and Appendix B in [11] and Section A in a Supplementary Material for examples.
This section studies generalized residual estimators $R^{f}_{k,n}$ (4.7), which are intrinsically related to generalized EKM integrals (5.1). The latter integrals are the main object of interest in Section 5, where the methodology developed in Section 3 is extended to all max-domains of attraction. The critical link is viewing generalized residual estimators as integrals over an empirical cdf of normalized order statistics of i.i.d. Pareto($1$) random variables. Upon replacement of this empirical cdf with the EKM estimator (5.2), we obtain a generalized EKM integral (5.1) capable of dealing with censoring effects. In fact, the same kind of replacement is implicitly made in Section 3, where standard EKM integrals could be seen as being derived from standard residual estimators.
As before, let $X_{1},\dots,X_{n}$ denote an i.i.d. sequence of random variables with cdf $F,$ and $X_{1,n}\leq\ldots\leq X_{n,n}$ denote the order statistics of $\{X_{i}\}_{i=1}^{n}.$ Denote by $U_{F}(t),$ the tail quantile function of $F,$ i.e. $U_{F}(t)=\inf\{x:F(x)\geq 1-1/t\}.$ When $F$ belongs to the Fréchet max-domain of attraction (that is $F\in\mathcal{D}(G_{\gamma})$ with $\gamma>0$), [28] studied the limit behavior of (standard) residual estimators defined by
$R_{k,n}=\frac{1}{k}\sum_{i=1}^{k}\vartheta(X_{n-i+1,n}/X_{n-k,n}),$ (4.1)
where $\vartheta:[1,\infty)\to\amsmathbb{R}$ is some suitable function, which is here assumed to be continuous.
Theorem 4.1 (Theorem 4.1 in [28]).
@@ 4 Generalized residual estimators for non-censored data
In further theorems 4.2 and 4.5 ibid., it was proven that $\mu(\xi_{n-k,n})$ in (4.2) can be replaced with the non-random limit $\amsmathbb{E}[\vartheta(\xi^{\gamma})]$ under some additional conditions imposed on $F,$ $\vartheta$ and $k.$
Residual estimators are closely related to the EKM integrals studied in Section 3. Indeed, the former can be represented as integrals over an empirical cdf built by normalized order statistics $\{X_{n-i+1,n}/X_{n-k,n}\}_{i=1}^{k}$:
$R_{k,n}=\int\vartheta\mathop{}\!\mathrm{d}F_{k,n},$ (4.3)
where $F_{k,n}(x)=k^{-1}\sum_{i=1}^{k}I(X_{n-i+1,n}/X_{n-k,n}\leq x).$ Replacing $F_{k,n}$ with the EKM estimator $\amsmathbb{F}_{k,n}$ (3.2) in (4.3) yields an EKM integral (3.3). The remainder of this section studies the fully-observed case. We refer to Section 5 for the generalization to censored data.
The first difficulty one encounters is that the direct extension of Theorem 4.1 to the Gumbel and Weibull max-domains of attraction leads to a degenerate limit in (4.2) in almost all cases. Recall the notation $\tau_{F}=\inf\{x:F(x)=1\},$ the right endpoint of $F.$ Specifically, we have the following result, whose proof is analogous to the proof of Theorem 4.1, and thus omitted.
Lemma 4.2.
@@
A reason of a degenerate limit in (4.4) is that for $\gamma\leq 0$ and $\tau_{F}>0$ the ratio $U_{F}(xt)/U_{F}(t),$ closely related to normalized order statistics, tends to $1$ for every $x>0$ as $t\to\infty$ (which follows from techniques in [11]), see for details Theorem 4.3 and Corollary A.1 of the Supplementary Material.
It follows that to derive a non-degenerate limit, a faster than $\sqrt{k}$ normalizing sequence is required in (4.4). Consequently, a downside of using a residual estimators technique to develop extreme value estimators (both in presence of and without random censorship) is that we may be faced with situations where positive and negative values of the extreme value index are allowed, but the rate of convergence of the estimator inconveniently varies. Another disadvantage is that $\vartheta(1)$ does not depend on $\gamma<0.$ This makes it impossible to construct extreme value index estimators based solely on $R_{k,n}.$
For this reason, let us consider an alternative and general approach, through generalized residual estimators. As before, let $\xi,\xi_{1},\xi_{2},\ldots$ be independent Pareto$(1)$ random variables with $n$-th order statistics denoted by $\xi_{1,n}\leq\ldots\leq\xi_{n,n},$ such that $X_{i}=U_{F}(\xi_{i})$, $i\geq 1$.
@@
and there exists $T>0$, such that for $t\geq T$
@@
We call $f$ a candidate function. If $f(x,t)$ tends to $\infty$ as $x\to x_{0}$ for some $x_{0}$ which can be infinite, we assume that $g_{1},\,g_{2}$ tend to $\infty$ together with $f(x,t)$.
Example 1.
Let $F\in\mathcal{D}(G_{\gamma}).$ Examples of candidate functions include
$\frac{U_{F}(xt)}{U_{F}(t)},\;\frac{U_{F}(xt)-U_{F}(t)}{a_{F}(t)},\;\frac{\log U_{F}(xt)-\log U_{F}(t)}{a_{F}(t)/U_{F}(t)},\quad\text{and}\quad\frac{\tau_{F}-U_{F}(xt)}{\tau_{F}-U_{F}(t)}$
with corresponding limits $x^{\gamma_{+}},$ $h_{\gamma}(x),$ $h_{\gamma_{-}}(x)$ and $x^{\gamma_{-}},$ as $t\to\infty,$ where $a_{F}(t)$ is a so-called auxiliary function, see, e.g., Theorem 1.1.6 in [11]. In Section A in Supplementary Material we discuss properties of the first three above candidate functions, whereas the first, third and fourth ones are utilized in Section 6.
Let $\vartheta:[0,\infty)\to\amsmathbb{R}$ be continuously differentiable and eventually monotone. We are interested in the limit behavior of the following generalized residual estimator:
$R^{f}_{k,n}=\frac{1}{k}\sum_{j=1}^{k}\vartheta\left(f\Big{(}\frac{\xi_{n-j+1,n}}{\xi_{n-k,n}},\xi_{n-k,n}\Big{)}\right)=\int\vartheta(f(x,\xi_{n-k,n}))F^{\xi}_{k,n}(\mathop{}\!\mathrm{d}x),$ (4.7)
where $k<n$ and $F^{\xi}_{k,n}(x)=k^{-1}\sum_{j=1}^{k}I(\xi_{n-j+1,n}/\xi_{n-k,n}\leq x).$ Note that taking $f(x,t)=U_{F}(xt)/U_{F}(t)$ in (4.7) we get the standard residual estimator (4.1).
@@
Let $k=k_{n}$ be an intermediate sequence and $\vartheta$ be continuous and eventually monotone. Assume for some $\delta>0$
$\amsmathbb{E}\Big{[}\max\big{\{}|\vartheta(g_{1}(\xi))|,|\vartheta(g_{2}(\xi))|\big{\}}\Big{]}^{1+\delta}<\infty.$ (4.9)
Then $R^{f}_{k,n}$ is consistent for $\amsmathbb{E}[\vartheta(g(\xi))].$ If, moreover, (4.9) is satisfied for some $\delta>1,$ then
@@
To replace $\mu_{f}(\xi_{n-k,n})$ with $\mu_{f}(\infty):=\amsmathbb{E}[\vartheta(g(\xi))]$ in Theorem 4.3, we need to impose another condition (compare with Condition 2).
Condition 3.
@@
Notice that all candidate functions from Example 1 satisfy Condition 3. For instance, for the candidate function $a^{-1}_{F}(t)U_{F}(t)(\log U_{F}(tx)-\log U_{F}(t))$ it reads as follows:
$\lim_{t\to\infty}\frac{\frac{\log U_{F}(tx)-\log U_{F}(t)}{a_{F}(t)/U_{F}(t)}-h_{\gamma_{-}}(x)}{Q(t)}=\int_{1}^{x}s^{\gamma_{-}-1}\int_{1}^{s}u^{\rho^{\prime}-1}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}s=:H_{\gamma_{-},\rho^{\prime}}(x),$ (4.13)
which for the cases $\gamma=0$ and $\rho=0$ is understood to be equal to the limit of (4.14) as $\gamma\to 0$ or $\rho\to 0,$ respectively.
@@
Let the conditions of Theorem 4.3 and Condition 3 hold. Assume $\sqrt{k}\hat{a}\big{(}(1-F(X_{n-k,n}))^{-1}\big{)}\stackrel{{\scriptstyle P}}{{\to}}\lambda$ with finite $\lambda,$ and $\vartheta^{\prime}$ is continuous and eventually monotone. Assume also
$\amsmathbb{E}\Big{[}\hat{g}(\xi)\max\big{\{}|\vartheta^{\prime}(g_{1}(\xi))|,|\vartheta^{\prime}(g_{2}(\xi))|\big{\}}\Big{]}<\infty.$ (4.15)